OpenStudy (anonymous):

State the domain of f(x)=log_3(x+1)- (x^2-4)^1/2

Parth (parthkohli):

Let's analyse both terms. \(\log_3(x+1)\) has a domain simply where \(x + 1 > 0\), or \(x > -1\). Similarly, \((x^2-4)^{1/2}\) has a domain where \(x^2 - 4 \ge 0\) or \((-\infty, -2] \cup [2, \infty)\). The above two conditions have to be simultaneously fulfilled.

OpenStudy (anonymous):

\[State the Domain for f(x)=\log_3{x+1} -\sqrt{x^2-4}\]

OpenStudy (anonymous):

Ok, Thank You very much! :)

Parth (parthkohli):

No problem! What do you get for the answer?

OpenStudy (anonymous):

\[(-\infty, -2] U [2,\infty) \]

OpenStudy (anonymous):

Since they both have to meet the requirement, then would that be correct?

Parth (parthkohli):

Remember that \(x > -1\) also, from the log term.

Parth (parthkohli):

Both requirements meaning: First: \(x \in (-1, +\infty)\) Second: \(x \in (-\infty,-2] \cup [2, \infty)\) Take the "intersection" of both.